# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ III-C Resource-Aware Model Optimization
Figure 5: Resource allocation strategies in edge-cloud system, which illustrates three approaches: (1) Dynamic resource allocation, (2) collaborative resource management, and (3) reinforcement learning optimization.
Dynamic adaptation to available resources and operational constraints emerges as a critical factor for model optimization in edge-cloud systems. As illustrated in Fig. 5, resource allocation integrates three complementary strategies to enhance ECCC performance across diverse environments. Dynamic resource allocation mechanisms continuously adapt to fluctuating workload demands and resource availability, orchestrating real-time service placement and task offloading decisions to maximize system efficiency. Building on this foundation, collaborative resource management techniques orchestrate allocation across multiple edge nodes and cloud resources, employing federated learning and hybrid expansion strategies to ensure cohesive operation. Further advancing this framework, reinforcement learning-based optimization refines system performance by integrating data-driven decisions informed by feedback on metrics such as latency and energy consumption, utilizing sophisticated methods like deep Q-networks and multi-agent coordination. Collectively, these strategies establish a versatile framework applicable to diverse domains, including IoT networks, drone communications, vehicular edge computing, and 5G environments.
III-C1 Dynamic Model Selection
Edge-cloud systems leverage distributed computing to optimize resource allocation and enhance application performance across environments. In ECCC, optimizing resource utilization and achieving desired performance necessitates selecting the most suitable AI model from a pool of candidates. Such dynamic model selection (DMS) process considers factors such as available resources at the edge and cloud (computation, communication bandwidth, and energy), task complexity, and desired QoS metrics, including latency and accuracy. Resource-constrained edge devices may require smaller, less computationally intensive models, potentially sacrificing some accuracy, while ample resources allow deploying more complex models for maximized performance [98]. Stringent latency requirements may prioritize models optimized for fast inference [53]. Additionally, DMS must adapt to the dynamic nature of edge-cloud environments, accommodating fluctuations in network connectivity or available resources to maintain optimal performance and resource efficiency [99]. Several research efforts have explored DMS strategies. Some leverage online learning techniques for dynamic adjustments based on real-time feedback [19]. Others employ reinforcement learning (RL) to learn optimal model selection policies for maximizing long-term performance [100]. Trust-aware DMS strategies incorporate the trustworthiness of edge resources into selection decisions [101].
III-C2 Model Splitting and Offloading
In ECCC, a crucial optimization technique involves partitioning a complex AI model into smaller sub-models and strategically deploying them across the edge-cloud continuum [20]. This distributed execution paradigm effectively leverages the heterogeneous nature of these systems, optimizing resource utilization and minimizing latency [102]. Computationally intensive components can be offloaded to the resource-rich cloud, while less demanding or latency-sensitive parts can be executed at the edge, closer to the data source [102]. For example, within a digital twin enabled 6G industrial IoT environment, this approach allows for efficient management of computationally intensive and delay-sensitive tasks within an end-edge-cloud architecture [103].
Several factors influence the optimization of model splitting and offloading decisions, including the computational capabilities of edge devices, communication bandwidth limitations, and application-specific latency requirements [104]. Resource-constrained edge devices often benefit from offloading complex layers of a CNN to the cloud [105]. Dynamic model selection and adaptation techniques, along with algorithms like multi-agent deep deterministic policy gradient (MADDPG), further enhance efficiency by tailoring the model architecture and execution strategy to the current system state and workload demands [106, 20]. Additionally, user mobility introduces further complexity, especially in heterogeneous edge-cloud computing scenarios with distributed resources [102].
III-C3 Reinforcement Learning for Model Optimization
In dynamic edge-cloud systems, RL provides a powerful mechanism for optimizing model parameters and offloading decisions [107]. RL agents learn optimal policies by leveraging environmental feedback, adapting model configurations and resource allocation strategies to improve system adaptability and performance. Specifically, techniques like deep Q-learning [108] and asynchronous proximal policy Optimization (APPO) [109] guide agents to explore actions and learn strategies that maximize cumulative rewards. For example, a study in [18] combined deep reinforcement learning with edge computing to optimize real-time monitoring and control in IoT environments, ultimately improving system stability and reducing failure rates. Additionally, RL enables sophisticated offloading mechanisms that consider factors such as latency, energy consumption, and task dependencies [110]. Through dynamic parameter adjustments, such as the number of layers or computational precision, RL optimizes the trade-off between model accuracy and computational cost [111]. Similarly, RL-based task offloading strategies dynamically determine optimal execution locations (edge or cloud) by adapting to real-time network and system conditions [109], simultaneously maximizing resource efficiency and reducing latency for time-critical applications.
@@ III-D Neural Architecture Search
For automating the design of optimal neural network architectures in edge-cloud systems, NAS methods aim to enhance both model performance and efficiency. NAS techniques are broadly categorized into three primary approaches: reinforcement learning-based, evolutionary algorithm-based, and gradient-based. In reinforcement learning-based NAS, an agent learns to select optimal architectures based on performance feedback. Evolutionary algorithm-based NAS, conversely, employs evolutionary algorithms to evolve populations of network architectures towards optimal solutions. Gradient-based NAS leverages gradient descent to directly search the architecture space for optimal configurations.
III-D1 Reinforcement Learning-based NAS
RL-NAS employs RL algorithms to automate the neural architecture search process. Specifically, an RL agent interacts with an environment representing the architecture search space and receives rewards based on the generated architectures’ performance, typically measured by validation set accuracy while considering edge-cloud deployment resource constraints [112]. Through iterative exploration and exploitation, the agent learns to select actions that result in high-performance architectures tailored to edge-cloud systems. [111]. Consequently, this approach can discover architectures that outperform manually designed ones, especially in resource-constrained edge devices [113]. For example, an RL agent can be trained to optimize for both accuracy and latency, balancing these conflicting objectives [22]. By incorporating resource constraints into the reward function, RL-NAS can directly target architectures suitable for edge-cloud environments with limited resources like computational power, memory, and energy budgets [114]. This adaptability to diverse edge environments makes RL-NAS promising for optimizing model deployment in ECCC [74].
III-D2 Evolutionary Algorithm-based NAS
For ECCC, neural architecture search (NAS) can leverage evolutionary algorithms (EAs) to discover optimal architectures by mimicking natural selection [115]. EAs encode network architectures as individuals within a population, subjecting them to iterative selection, mutation, and crossover to enhance performance [116]. Selection favors architectures demonstrating superior accuracy, latency, or energy efficiency [117], while mutation introduces random architectural changes to explore new design possibilities [118]. Crossover combines elements of high-performance architectures, potentially yielding even better solutions [19]. This iterative refinement identifies architectures suited for the resource constraints and performance needs of edge-cloud environments [116]. For instance, the self-adaptive discrete particle swarm optimization algorithm with genetic algorithm operators (GA-DPSO) optimizes data transmission time for scientific workflows in edge-cloud systems [119]. The inherent adaptability of EAs proves particularly valuable in dynamic edge-cloud environments characterized by fluctuating resource availability and workload demands [120].
III-D3 Gradient-based NAS
Leveraging gradient descent, gradient-based NAS methods efficiently explore and optimize the architecture search space [19]. Unlike earlier discrete search methods, this approach formulates architecture search as a continuous optimization problem, enabling the use of established gradient-based optimization techniques [121]. Specifically, the architecture search space is represented differentiably, allowing gradient descent to update architectural parameters and iteratively refine the model for optimal performance and efficiency. Consequently, search time is significantly reduced compared to discrete search methods, facilitating exploration of complex architecture spaces and tailoring models for diverse edge-cloud environments [113]. For example, DARTS [121] constructs a differentiable search space and optimizes it using gradient descent, achieving high-performance architectures rapidly. Despite these advantages, challenges remain, such as handling strong correlations among operations, which can lead to suboptimal architectures. In addition, techniques like weight sharing and channel bottlenecks can further enhance efficiency by reducing the computational cost of evaluating different architectures [122]. Critically, incorporating device constraints, including model size and floating-point operations, directly into the search process is essential for deploying optimized models on resource-constrained edge devices [122].
@@ IV-A Task Offloading Mechanisms
TABLE IV: Summary of task offloading mechanisms optimizing latency, energy efficiency, and dependency management.
Method Domain Performance Metrics Performance Results Edge Cloud ECC Task Off.
Zhang et al. [19] Edge Cloud Architecture Time, Cost, Load balance, User satisfaction, Trust measurement Found more than half of the best solutions on benchmark functions
Fan et al. [123] Edge-Cloud Cooperative Computing Task processing delay, Task queuing stability Superior performance compared to 4 other schemes
Qi et al. [99] Edge-Cloud Service Provisioning Service quality, Cost optimization Effective cost-service quality balance demonstrated
Abouaomar et al. [124] Edge Computing for IoT Latency, Resource consumption Outperforms benchmark approaches in latency and resource utilization
Li et al. [125] Edge-Cloud Cooperation Long-term network performance, Energy constraints Converges to approximate optimal solution in polynomial time
Luckow et al. [126] Edge-to-Cloud Applications Performance impact of task placement 5-65% performance improvement depending on scenario
Cai et al. [127] Data-Intensive Edge Services Throughput, Delay, Resource consumption Superior performance compared to state-of-the-art algorithms
Pang et al. [128] Vehicle-Edge-Cloud Latency, Energy consumption Significant decrease in latency and energy consumption
PRAISE [129] Multi-Access Edge Cloud System benefits, Resource costs Outperforms baselines with higher benefits and lower costs
Khalafi et al. [130] Edge-Cloud Orchestrated Networks System cost, QoS, Delay violations Reduced costs associated with network slicing
Peng et al. [131] Edge Computing Systems Job computing time, Job blocking probability Optimal replication factor varies with distribution parameters and arrival rate
DRLIS [132] Edge-Fog-Cloud Computing Load balancing, Response time, Weighted cost Reduces execution cost by up to 55% (load balancing), 37% (response time), 50% (weighted cost)
Hua et al. [102] Heterogeneous Edge-Cloud Computing Energy consumption Significantly reduced energy consumption of mobile devices
Arroba et al. [133] Sustainable Edge Computing Energy efficiency, Economic viability Early results demonstrate benefits for future computing environments
PDCO [134] Cloud-Edge Collaboration Total energy consumption, Resource utilization Approaches optimal performance in energy consumption and resource utilization
He et al. [135] Wireless Powered MEC Systems Amount of processed data Enhances computation performance compared to benchmark methods
@@
iGATS [8] IoT-Edge-Cloud Networks Response latency, Temporal satisfaction, Load-balancing Outperforms baseline techniques for reducing response latency
@@
ETCRA [101] Edge-Cloud Workflows Comprehensive system function, Execution time, Energy consumption, Reliability Significantly outperforms other baselines on four key measurements
Hu et al. [136] MEC-enabled Mobile IoT Energy efficiency, Service delay Achieves [O(1/V), O(V)] tradeoff between EE and service delay
Alopex-DE [116] Local-Edge-Cloud Computing Latency, Energy consumption Superior effectiveness compared to three state-of-the-art algorithms
@@
DT-MADDPG [103] Digital Twin Enabled 6G IIoT Task success rate, Latency, Energy consumption Significantly improved task success rate with reduced latency and energy consumption
Fan et al. [137] Cloud-Edge-End Cooperative Networks Task processing delay, Energy consumption, Block mining speed Superior performance compared to 5 reference schemes in 7 scenarios
@@
Zhang et al. [138] Edge-Cloud Systems Task completion time, Cost, Energy consumption, System reliability Outperforms existing design approaches
Qi et al. [99] Edge-Cloud Collaboration Cost, Service quality Effective in both synthetic and real-world data experiments
@@
EdgeOptimizer [139] Kubernetes-Based Edge-Cloud Clusters Success rate of time-critical tasks, Overall overload Improved success rate for time-critical tasks
@@
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 6, a framework outlines three core strategies—latency-aware, energy-efficient, and dependency-aware offloading—detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table IV summarizes recent task offloading methods in edge-cloud environments. Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [140], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques. In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [136]. As shown in Table IV, numerous approaches including PDCO [134], Hua et al. [102], and ETCRA [101] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [136, 135]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [102], and user mobility patterns [133]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [116], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [140]. Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [102]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [133]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [116]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [140].
@@ IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [129]. Such interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [129] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [128]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [141], enabling dynamic adaptation to changing environments without pre-set task priorities.

### v4

@@ III-C Resource-Aware Model Optimization
Resource-aware model optimization represents a critical advancement in ECCC systems, focusing on dynamic adaptation to heterogeneous and constrained computational environments. Advanced optimization strategies encompass three primary approaches: dynamic model selection based on real-time resource availability and application requirements, model splitting and offloading that strategically partition components across edge-cloud infrastructure, and reinforcement learning-based optimization for adaptive resource allocation. Dynamic model selection considers computational constraints, communication bandwidth, and energy limitations while maintaining QoS requirements, with techniquess ranging from online learning approaches to trust-aware selection strategies [116, 62]. Model partitioning leverages heterogeneous system capabilities by offloading computationally intensive components to resource-rich cloud while executing latency-sensitive parts at the edge, with sophisticated algorithms like MADDPG enhancing efficiency through workload-aware adaptation [23, 117]. Reinforcement learning mechanisms optimize resource allocation through environmental feedback, employing techniques like deep Q-learning and APPO to balance computational efficiency with system performance, achieving improved stability and reduced failure rates in IoT monitoring applications [21, 118]. Consequently, integrated optimization strategies establish a versatile framework for diverse domains including IoT networks, vehicular computing, and 5G environments, addressing the fundamental challenges of resource heterogeneity and dynamic operational constraints in distributed computing systems. For comprehensive technical details, evaluation methodologies, and implementation strategies, please see LABEL:sec:supp-tech-comparison of the Supplementary Material.
@@ III-D Neural Architecture Search
Neural architecture search (NAS) automates the design and optimization of neural network architectures to discover optimal configurations for edge-cloud collaborative environments, systematically exploring architectural configurations tailored to distributed computing constraints. Advanced automated methodologies encompass three primary paradigms: reinforcement learning-based NAS where agents learn optimal architecture selection through performance feedback and reward systems that incorporate edge-cloud deployment constraints; evolutionary algorithm-based NAS that employs natural selection principles to evolve architecture populations through iterative selection, mutation, and crossover operations; and gradient-based NAS that formulates architecture search as continuous optimization problems enabling efficient exploration through established gradient descent techniques [119, 120, 22]. Reinforcement learning approaches demonstrate particular effectiveness in resource-constrained environments by directly incorporating computational, memory, and energy constraints into reward functions, enabling discovery of architectures specifically optimized for edge device limitations [25, 121]. Evolutionary algorithms provide exceptional adaptability to dynamic edge-cloud environments through inherent population diversity and exploration capabilities, with applications ranging from scientific workflow optimization to adaptive resource allocation [122, 123]. Gradient-based methods achieve superior search efficiency by representing architecture spaces differentiably, reducing search time significantly while enabling exploration of complex architectural configurations, though challenges remain in handling operation correlations and incorporating device constraints directly into the search process [124, 125]. These complementary approaches collectively enable automated discovery of neural architectures that outperform manually designed alternatives while meeting the stringent performance and resource requirements of distributed edge-cloud deployments. For detailed technical analysis, comparative evaluation, and implementation guidelines, please see LABEL:sec:supp-resource-nas of the Supplementary Material.
@@ IV AI-Driven Resource Management
Effective resource management in ECCC systems requires intelligent orchestration of computational, communication, and storage resources across heterogeneous infrastructure to optimize system performance while meeting diverse application requirements. The section examines three complementary yet distinct aspects of AI-driven resource management. Task offloading mechanisms (Sec. IV-A) focus on intelligent decision-making for task placement across the edge-cloud continuum, considering factors such as latency requirements, energy constraints, and task dependencies to determine optimal execution locations. Resource allocation strategies (Sec. IV-B) address the dynamic distribution and management of computational resources (CPU, memory, storage) across edge nodes and cloud servers, emphasizing adaptive allocation algorithms and collaborative resource sharing approaches. Energy efficiency optimization (Sec. IV-C) concentrates on minimizing power consumption throughout the system lifecycle, including power modeling, energy-aware scheduling, and joint optimization techniques that balance performance with sustainability requirements.
@@ IV-A Task Offloading Mechanisms
TABLE IV: Summary of task offloading mechanisms, categorized by strategy type and performance optimization targets.
Type Method Metrics Achievements Gain ECC
Latency-Aware
@@
iGATS [8] Response latency, Load balancing Outperforms baseline techniques for latency reduction High
Fan et al. [126] Task processing delay, Stability Superior performance vs. 4 comparison schemes Medium
Zhai et al. [127] Response delay, Carbon emissions 33.42× delay reduction for sensitive workloads Very High
Luckow et al. [128] Performance impact, Task placement 5-65% performance improvement across scenarios High
Energy-Efficient
PDCO [129] Energy consumption, Resource util. Approaches optimal performance in energy metrics High
Hua et al. [117] Energy consumption of mobile devices Significantly reduced energy consumption High
@@
ETCRA [130] Energy, Time, Reliability Outperforms baselines on four key measurements High
Hu et al. [131] Energy efficiency, Service delay Achieves [O(1/V), O(V)] EE-delay tradeoff Medium
Dependency-
Aware
PRAISE [132] System benefits, Resource costs Higher benefits and lower costs vs. baselines Medium
Peng et al. [133] Computing time, Blocking probability Optimal replication with varying parameters Medium
Multi-Objective
Zhang et al. [22] Time, Cost, Load balance, Trust Found >>50% of best solutions on benchmarks High
DRLIS [134] Load balancing, Response time, Cost Up to 55% cost reduction across metrics Very High
@@
DT-MADDPG [135] Success rate, Latency, Energy Improved success rate with reduced resource usage High
@@
Zhang et al. [136] Time, Cost, Energy, Reliability Outperforms existing design approaches Medium
@@
EdgeOptimizer [137] Success rate, Overall overload Improved success rate for time-critical tasks Medium
Notes: Performance Gain: Very High (>>50%), High (20-50%), Medium (5-20%).
@@
Effective resource utilization and performance optimization in ECCC depend critically on sophisticated task offloading mechanisms. As depicted in Fig. 5, a framework outlines three core strategies: latency-aware, energy-efficient, and dependency-aware offloading, detailing their interactions with edge devices, cloud servers, and supporting systems to improve task placement and scheduling. Specifically, these mechanisms address distinct objectives such as real-time performance, energy consumption, and task dependencies, thereby boosting overall system efficiency. To further illustrate their practical impact, Table IV summarizes recent task offloading methods in edge-cloud environments.
Notably, the table reveals significant performance gains, such as up to a 33.42× reduction in delay for sensitive workloads [127], alongside improvements in energy consumption and system reliability, highlighting the versatility of these strategies across diverse application domains.
@@ IV-A2 Energy-Efficient Offloading
Strategic energy management across edge-cloud systems requires balancing computation and communication costs while optimizing resource allocation through specialized techniques.
In ECCC, prolonging the battery life of edge devices necessitates balancing computation and communication energy costs, especially for battery-powered devices [131]. As shown in Table IV, numerous approaches including PDCO [129], Hua et al. [117], and ETCRA [130] have demonstrated significant improvements in energy efficiency across edge-cloud systems. Particularly for battery-powered devices, strategic decision-making has to consider both computational and communication energy costs, weighing factors like task complexity, network conditions, and hardware energy profiles [131, 141]. While offloading computationally intensive tasks to cloud resources can reduce device energy consumption, communication costs under limited bandwidth or unstable networks often counterbalance these benefits, necessitating sophisticated optimization approaches. Beyond basic consumption metrics, advanced energy-efficient mechanisms consider additional dimensions such as power control, transmission scheduling [117], and user mobility patterns [142]. Furthermore, in heterogeneous IoT deployments with diverse device capabilities and constraints, efficient resource allocation becomes increasingly crucial for sustainable service delivery [122], with some approaches even incorporating renewable energy sources to create more environmentally responsible edge-cloud ecosystems [127].
Energy-efficient offloading mechanisms strive to minimize overall energy consumption by jointly considering computation and communication costs. Additionally, optimizing power control and transmission scheduling can further enhance energy efficiency [117]. User mobility introduces another layer of complexity to energy optimization, as dynamic network conditions and varying proximity to edge servers impact energy trade-offs [142]. In heterogeneous IoT environments with diverse device capabilities and energy constraints, efficient resource allocation becomes even more crucial for energy-efficient service delivery [122]. Prioritizing workloads based on their energy requirements and leveraging renewable energy sources can contribute to a more sustainable edge-cloud ecosystem [127].
@@ IV-A3 Dependency-Aware Offloading
Complex applications often consist of multiple interconnected subtasks with dependencies, forming a dependency graph where some subtasks must finish before others can start [132]. Complex interdependence introduces complexities in task scheduling and resource allocation, potentially leading to suboptimal performance due to idle resources and increased latency if ignored. Therefore, efficient offloading mechanisms must consider these precedence constraints to optimize overall application execution. For example, the PRAISE algorithm in [132] addresses triangular dependencies by reconfiguring the task call graph into sequential layers for efficient resource allocation. Similarly, modeling task dependencies as a directed acyclic graph (DAG) optimizes offloading decisions, particularly in vehicle-edge-cloud environments [139]. Additionally, deep Q-networks (DQN) for real-time offloading in cloud-edge computing benefit from incorporating dependency awareness [143], enabling dynamic adaptation to changing environments without pre-set task priorities.
IV-A4 Multi-Objective Optimization Strategies
Contemporary edge-cloud systems increasingly require sophisticated offloading strategies that simultaneously optimize multiple performance metrics, reflecting the complex trade-offs inherent in distributed computing environments. Unlike single-objective approaches that focus on isolated metrics, multi-objective optimization techniques explicitly address the conflicts between competing goals such as minimizing latency while reducing energy consumption, maximizing throughput while maintaining quality of service, and balancing computational load while ensuring system reliability [22]. Multi-objective approaches typically employ Pareto optimization methods to identify solution sets that represent optimal trade-offs between conflicting objectives, enabling system designers to select configurations that best match their specific application requirements and operational constraints [144].
Recent research demonstrates the effectiveness of multi-objective approaches across diverse scenarios. For instance, Zhang et al. [22] propose a many-objective resource scheduling model that simultaneously considers time, cost, load balance, user satisfaction, and trust factors, employing advanced evolutionary algorithms to navigate the complex solution space. Similarly, Guo et al. [145] frame service deployment and task offloading as a multi-task, high-dimensional optimization problem that incorporates network reliability and load balancing considerations alongside traditional performance metrics. Furthermore, joint optimization techniques that combine energy efficiency with latency minimization have proven particularly effective in battery-powered edge environments [117], while approaches that integrate security and privacy considerations with performance optimization address the growing demands of trust-sensitive applications [130]. The challenge lies in developing efficient algorithms that can handle the increased complexity of multi-dimensional solution spaces while maintaining real-time decision-making capabilities essential for dynamic edge-cloud environments.